OpenStudy (anonymous):

Felix purchased a boat for $9,755. He made a down payment of $875. He applied for a six-year installment loan with an interest rate of 10.4% in the amount of $8,880. What is the total cost of the boat after six years? $9,803.52 $11,974.32 $12,849.32 $10,769.52 Need some major help understanding this! Please Help!

OpenStudy (math&ing001):

The bank is taking an interest rate of 10.4% from the loan you took wich is $8880. So the total cost would be the amount $8880 plus the interest the bank is taking ie (10.4*8880)/100 Total cost = 8880+0.104*8880 = $9803.52

OpenStudy (anonymous):

Thank you I thought it was that answer but its wrong

OpenStudy (math&ing001):

Yeah sry just googled it, apparently installment loan is different than the normal one. cf http://en.wikipedia.org/wiki/Installment_loan

OpenStudy (anonymous):

Yes its okay you have to use this formula M= P ( i(+1)^nt / (1+i)^nt -1) but I keep doing something wrong

OpenStudy (anonymous):

I got the answer is 12849

OpenStudy (math&ing001):

Gratz !